### v1

# Solvability of the Korteweg-de Vries equation under meromorphic initial conditions by quadrature

Abstract.
We study the solvability of the Korteweg-de Vries equation under meromorphic initial conditions by quadrature when the inverse scattering transform (IST) is applied. It is a key to solve the Schrödinger equation appearing in the Lax pair in application of the IST. We show that the Schrödinger equation is always integrable in the sense of differential Galois theory, i.e., solvable by quadrature, if and only if the meromporphic potential is reflectionless, under the condition that the potential is absolutely integrable on $\mathbb{R}\setminus(-R_{0},R_{0})$ for some $R_{0}>0$. This statement was previously proved to be true by the author for a limited class of potentials. We also show that the Schrödinger equation is not integrable in this sense for rational potentials that decay at infinity but do not satisfy the weak condition.
Key words and phrases:
Solvability by quadrature; Korteweg-de Vries equation; Schrödinger equation; inverse scattering transform; differential Galois theory; meromorphic initial condition; reflectionless potential
2020 Mathematics Subject Classification:
35Q53, 37K15, 34M03, 34M15, 34M35, 34M40, 35P25
1. Introduction
In this paper we study the solvability of the Korteweg-de Vries (KdV) equation
$u_{t}+6uu_{x}+u_{xxx}=0$ (1.1)
under meromorphic initial conditions by quadrature, where the subscripts represent differentiation with respect to the independent variables. Here the solution to (1.1) are required to satisfy
$\lim_{x\to\pm\infty}u(x,t)=0.$
We also pay special attention to the case in which the initial conditions are given by rational functions.
As well-known, the initial value problems (IVPs) of the KdV equation (1.1) can be solved by the inverse scattering transform (IST) (see, e.g., Chapter 9 of [1]). In the method, a pair of differential equations
$\varphi_{xx}+u(x,t)\varphi=\lambda\varphi,$ (1.2)
$\varphi_{t}=-4\varphi_{xxx}-6u(x,t)\varphi_{x}+(\mu-3u_{x}(x,t))\varphi,$ (1.3)
which are called the Lax pair, are analyzed to obtain the solutions to the IVP, where $\lambda,\mu\in\mathbb{C}$ are constants. In particular, we need to solve (1.2) for any $\lambda\in\mathbb{C}^{\ast}=\mathbb{C}\setminus\{0\}$ and obtain particular solutions called the Jost solutions, which satisfy
$\begin{split}&\phi(x;k)\sim e^{-ikx}\quad\mbox{as $x\to-\infty$},\\ &\psi(x;k)\sim e^{ikx}\quad\mbox{as $x\to+\infty$},\end{split}$ (1.4)
where $\lambda=-k^{2}\neq 0$. Thus, when the IST is applied, the IVP of (1.1) is solvable by quadrature only if so is (1.2) for any $\lambda\in\mathbb{C}$. We discuss the solvability of (1.1) by quadrature in this meaning, reducing it to that of the Schrödinger equation
$v_{xx}+(k^{2}+u(x))v=0,$ (1.5)
or as a first-order system
$w_{x}=\begin{pmatrix}0&1\\ -k^{2}-u(x)&0\end{pmatrix}w,\quad w=\begin{pmatrix}v\\ \mathrm{d}v/\mathrm{d}x\end{pmatrix}\in\mathbb{C}^{2},$ (1.6)
i.e., its integrability in the sense of differential Galois theory [10, 22]. See also Section 3 of [26] for a quick review of a necessary part of the theory.
Such integrability of (1.5) (and equivalently (1.6)) was recently discussed in [26] under the following condition:
  1. (A0)
    The potential $u(x)$ is holomorphic in a neighborhood $U$ of $\mathbb{R}$ in $\mathbb{C}$. Moreover, there exist holomorphic functions $u_{\pm}:U_{0}\to\mathbb{C}$ such that $u_{\pm}(0)=0$ and
    $u(x)=u_{\pm}(e^{\mp\lambda_{\pm}x})$
    for $|\mathrm{Re}\,x|$ sufficiently large, where $U_{0}$ is a neighborhood of the origin in $\mathbb{C}$, $\lambda_{\pm}\in\mathbb{C}$ are some constants with $\mathrm{Re}\,\lambda_{\pm}>0$, and the upper or lower signs are taken simultaneously depending on whether $\mathrm{Re}\,x>0$ or $\mathrm{Re}\,x<0$.
In particular, $u(x)$ tends to zero exponentially as $x\to\pm\infty$ on $\mathbb{R}$, so that $u\in L^{1}(\mathbb{R})$, if it satisfies condition (A0). Condition (A0) is a little restrictive, but it is satisfied by several wide classes of functions. For example, if $u(x)$ is a rational function of $e^{\lambda x}$ for some $\lambda\in\mathbb{C}$ with $\mathrm{Re}\,\lambda>0$, has no singularity on $\mathbb{R}$, and $u(x)\to 0$ as $x\to\pm\infty$, then condition (A0) holds. See Section 1 of [26] for another example of $u(x)$ satisfying condition (A0).
Define the scattering coefficients $a(k)$ and $b(k)$ for (1.5) as
$\phi(x;k)=a(k)\psi(x;-k)+b(k)\psi(x;k).$ (1.7)
When $a(k)\neq 0$, the constant $\rho(k)=b(k)/a(k)$ is defined and called the reflection coefficient. The potential $u(x)$ is called reflectionless if $b(k)=0$ for any $k\in\mathbb{R}^{\ast}:=\mathbb{R}\setminus\{0\}$. The following result was proved in [26].
Theorem 1.1.
Suppose that condition (A0) holds. If the potential $u(x)$ is reflectionless, then Eq. (1.5) is solvable by quadrature for any $k\in\mathbb{C}^{\ast}$. Conversely, if Eq. (1.5) is solvable by quadrature for any $k\in\mathbb{R}^{\ast}$, then the potential $u(x)$ is reflectionless.
Remark 1.2.
A similar statement was proved for more general two-dimensional systems called the Zakharov-Shabat ((((ZS)))) systems, which appear in application of the IST for other integrable partial differential equations ((((PDEs)))) including the nonlinear Schrödinger equation, modified KdV equation, sine-Gordon equation and sinh-Gordon equation, in [26].
In this paper we consider the case in which $u(x)$ is meromorphic, and extend the result of [26]. More precisely, we assume the following.
  1. (A1)
    The potential $u(x)$ is meromorphic in a neighborhood of $\mathbb{R}$ in $\mathbb{C}$ and
    $\int_{\mathbb{R}\setminus(-R_{0},R_{0})}|u(x)|\mathrm{d}x<\infty\quad\mbox{forsome $R_{0}>0$.}$ (1.8)
When Eq. (1.8) holds, we have
$\lim_{x\to\pm\infty}u(x)=0,$ (1.9)
and by Theorem 8.1 in Section 3.8 of [9] there exist the Jost solutions $\phi(x;k),\psi(x;k)$ satisfying (1.4) for $k\in\mathbb{C}^{\ast}$. See also Section 2 of [26]. On the other hand, meromorphic functions that are not analytic have singular points. Singular solitons called positons, negatons and complexitons in the KdV equation (1.1) have attracted much attention and been studied (see, e.g., [14, 16, 17, 18, 19, 24]). Moreover, rational solitons, which have singularities, in the KdV equation (1.1) were studied in [4, 5, 8, 11]. In particular, Jiménez et al. [11] discussed the integrability of the Schrödinger equation (1.5) in the sense of differential Galois theory for the Adler-Moser rational potentials [4, 5], for example,
$u(x)=\frac{2}{x^{2}},\ \frac{6x(x^{3}-6\tau)}{(x^{3}+3\tau)^{2}},\ \frac{6x(2x^{9}+675\tau^{2}x^{3}+1350\tau^{3})}{(x^{6}+15\tau x^{3}-45\tau^{2})^{2}},\ \ldots,$ (1.10)
where $\tau\in\mathbb{C}$ is any constant, which correspond to rational solitons in the KdV equation (1.1). See [4, 11] for more details of the Adler-Moser rational potentials including their precise definition. They showed that its differential Galois group is trivial for $k=0$ and diagonalizable for $k\neq 0$. This result implies that Eq. (1.5) is solvable by quadrature for the Adler-Moser rational potentials. We also notice that Acosta-Humánez et al. [3] and Blázquez-Sanz and Yagasaki [7] studied the integrability of the Schrödinger equation (1.5) for rational potentials and potentials satisfying condition (A0), respectively, in the sense of differential Galois theory.
Let ±=$\mathbb{C}_{\pm}=\{k\in\mathbb{C}\mid\pm\mathrm{Im}\,k>0\}$. Our main results are stated as follows.
Theorem 1.3.
Suppose that condition (A1) holds and $a(k)$ has a zero in $\mathbb{C}_{+}$. If $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$, then Eq. (1.5) is solvable by quadrature for any $k\in\mathbb{C}^{\ast}$.
Theorem 1.4.
Suppose that condition (A1) holds and $u(x)$ is analytic in a neighborhood of $x=\infty$ in the Riemann sphere ${}\mathbb{C}\cup\{\infty\}$. If Eq. (1.5) is solvable by quadrature for any $k\in\mathbb{C}^{\ast}$, then $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$.
Remark 1.5.
The scattering coefficient $a(k)$ may have no zero in $\mathbb{C}_{+}$ when condition (A1) holds and $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$. Actually, for the Adler-Moser rational potentials (1.10), condition (A1) holds but $a(k)=1$ and $b(k)=0$ for any $k\in\mathbb{C}^{\ast}$, as shown in [11]. In such a case, Theorem 1.3 does not apply.
Furthermore, we consider the case in which $u(x)$ is a rational function and satisfies (1.9) but does not (1.8).
Theorem 1.6.
Suppose that $u(x)$ is a rational function, and let $m_{1}$ and $m_{2}$ be the degrees of its numerator and denominator. If $m_{2}-m_{1}=1$, then Eq. (1.5) is not solvable by quadrature for some $k\in\mathbb{R}^{\ast}$.
Obviously, if $m_{2}-m_{1}>1$, then the rational function $u(x)$ satisfies (1.8) and condition (A1).
The outline of this paper is as follows. In Section 2 we collect preliminary results for proving Theorems 1.3 and 1.4. We give proofs of Theorems 1.3, 1.4 and 1.6 in Sections 3, 4 and 5, respectively. The result of Kovacic [13], which plays a crucial role in Section 5, is briefly reviewed in Appendix A.
2. Preliminary Results
In this section we give preliminary results required to prove Theorems 1.3 and 1.4. Henceforth we assume that condition (A1) holds.
Let $v_{l}(x)$, $l=1,2$, be scalar functions on $\mathbb{R}$ or $\mathbb{C}$, and let $W(v_{1}(x),v_{2}(x))$ denote the Wronskian of $v_{1}(x)$ and $v_{2}(x)$, i.e.,
$W(v_{1}(x),v_{2}(x))=\begin{vmatrix}v_{1}(x)&v_{2}(x)\\ v_{1x}(x)&v_{2x}(x)\end{vmatrix}=v_{1}(x)v_{2x}(x)-v_{2}(x)v_{1x}(x).$ (2.1)
Since the first-order derivative $v_{x}$ does not appear in (1.5), the Wronskians of $\phi(x;-k),\phi(x;k)$ and of $\psi(x;k),\psi(x;-k)$ are independent of $x$, so that by (1.4)
$W(\phi(x;k),\phi(x;-k))=W(\psi(x;-k),\psi(x;k))=2ik.$ (2.2)
Replacing $k$ with $-k$ in (1.7), we have
$\phi(x;-k)=a(-k)\psi(x;k)+b(-k)\psi(x;-k),$ (2.3)
and consequently
$a(k)a(-k)-b(k)b(-k)=1$ (2.4)
by (2.2). We have the following properties of the scattering coefficients.
Proposition 2.1.
  • (i)
    $a(k),b(k)$ are analytic in $\mathbb{R}^{\ast};$
  • (ii)
    $a(k)$ can be analytically continued in $\mathbb{C}_{+};$
  • (iii)
    $a(k)$ only has discrete zeros in $\mathbb{C}_{+}\cup\mathbb{R};$
  • (iv)
    If $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$, then $a(k)$ only has finitely many zeros in $\mathbb{C}_{+}$.
Similar results are obtained in Section 4.1 (especially Proposition 4.1) of [26] when $u(x)$ is analytic on $\mathbb{R}$.
Proof.
Let $S$ denote the set of poles of $u(x)$. Then $S$ is discrete and contains no accumulation point. Choose a curve $\Gamma=\{\gamma(\xi)\in\mathbb{C}_{+}\mid\xi\in\mathbb{R}\}$ with
$\gamma(\xi)=\xi+ic\operatorname{sech}\xi$
does not intersect $S$. Hence, $u(x)$ is analytic on $\Gamma$.
Regard (1.5) as a differential equation on $\Gamma$ and let $U$ be a neighborhood of $\Gamma$ in $\mathbb{C}_{+}$ such that $S\cap U=\emptyset$. Then its solutions $v=\phi(x;k)$ and $\psi(x;k)$ are bounded and analytic in $k\in\mathbb{R}^{\ast}$ as well as in $x\in U$. By (1.7) and (2.1) we have
$a(k)=\frac{W(\phi(x,k),\psi(x;k))}{2ik},\quad b(k)=-\frac{W(\phi(x,k),\psi(x;-k))}{2ik}.$ (2.5)
Hence, we obtain part (i). Similarly we see that $v=\phi(x;k)$ and $\psi(x;k)$ are bounded and analytic in $\mathbb{C}_{+}$ for $x\in U$ satisfying $|x|<R$ with some $R>0$. This yields part (ii) along with (2.5).
Equation (1.5) is written as
$v_{\xi\xi}-\frac{\gamma_{\xi\xi}(\xi)}{\gamma_{\xi}(\xi)}v_{\xi}+\gamma_{\xi}(\xi)(k^{2}+u(\gamma(\xi)))v=0$ (2.6)
on $\Gamma$. Changing the independent variable as $\xi\mapsto k\xi$ in (2.6), we have
$v_{\xi\xi}-\frac{\gamma_{\xi\xi}(\xi/k)}{k\gamma_{\xi}(\xi/k)}v_{\xi}+\gamma_{\xi}(\xi/k)\left(1+\frac{1}{k^{2}}u(\gamma(\xi/k))\right)v=0,$
which reduces to
$v_{\xi\xi}+v=0$
as $k\to\infty$. This implies that
$a(k)\to 1,\quad b(k)\to 0\quad\mbox{as $|k|\to\infty$}.$
We apply the identity theorem (e.g., Theorem 3.2.6 of [2]) to obtain part (iii), since $a(k)$ is analytic in $\mathbb{C}_{+}\cup\mathbb{R}$. If $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$, then by its analyticity $a(k)$ has no zero near $k=0$ in $\overline{\mathbb{C}}_{+}$. Hence, part (iv) follows from part (iii). ∎
Remark 2.2.
The conclusion in Proposition 2.1(iv) holds under a weaker condition. For example, if $\lim_{k\to 0}b(k)b(-k)\to b_{0}$ $(\neq-1)$, then it holds.
Let $G$ be the differential Galois group of (1.5). Then $G$ is an algebraic group such that $G\subset\mathrm{SL}(2,\mathbb{C})$ since Eq. (1.5) has no term depending on $v_{x}$ (see, e.g., Chapter 6 of [12] or Section 2.2 of [21]). We have the following classification for such algebraic groups (see Section 2.1 of [21] for a proof). Recall that an algebraic group $\mathscr{G}$ contains in general a unique maximal algebraic subgroup $\mathscr{G}^{0}$, which is called the connected component of the identity or connected identity component.
Proposition 2.3.
Any algebraic group $\mathscr{G}\subset\mathrm{SL}(2,\mathbb{C})$ is similar to one of the following types:
  1. (i)
    $\mathscr{G}$ is finite and $\mathscr{G}^{0}=\{\mathrm{id}\}$, where $\mathrm{id}$ is the $2\times 2$ identity matrix;
  2. (ii)
    $\mathscr{G}=\left\{\begin{pmatrix}\lambda&0\\ \mu&\lambda^{-1}\end{pmatrix}\middle|\,\lambda\text{ is a root of $1$, $\mu\in\mathbb{C}$}\right\}$ and $\mathscr{G}^{0}=\left\{\begin{pmatrix}1&0\\ \mu&1\end{pmatrix}\middle|\,\mu\in\mathbb{C}\right\}$;
  3. (iii)
    $\mathscr{G}=\mathscr{G}^{0}=\left\{\begin{pmatrix}\lambda&0\\ 0&\lambda^{-1}\end{pmatrix}\middle|\,\lambda\in\mathbb{C}^{*}\right\}$;
  4. (iv)
    $\mathscr{G}=\left\{\begin{pmatrix}\lambda&0\\ 0&\lambda^{-1}\end{pmatrix},\begin{pmatrix}0&-\beta^{-1}\\ \beta&0\end{pmatrix}\middle|\,\lambda,\beta\in\mathbb{C}^{*}\right\}$ and $\mathscr{G}^{0}=\left\{\begin{pmatrix}\lambda&0\\ 0&\lambda^{-1}\end{pmatrix}\middle|\,\lambda\in\mathbb{C}^{*}\right\}$;
  5. (v)
    $\mathscr{G}=\mathscr{G}^{0}=\left\{\begin{pmatrix}\lambda&0\\ \mu&\lambda^{-1}\end{pmatrix}\middle|\,\lambda\in\mathbb{C}^{*},\,\mu\in\mathbb{C}\right\}$;
  6. (vi)
    $\mathscr{G}=\mathscr{G}^{0}=\mathrm{SL}(2,\,\mathbb{C})$.
The system (1.5) is integrable in the sense of differential Galois theory unless $G$ is of type (vi). Proposition 2.3 plays a key role in the proof of Theorem 1.4 in Section 4.
3. Proof of Theorem 1.3
In this section we give a proof of Theorem 1.3. Henceforth we assume that the hypotheses of Theorem 1.3 hold and $b(k)=0$ for $k\in\mathbb{R}^{\ast}$.
Proof of Theorem 1.3.
We see by Proposition 2.1(iii) and (2.4) that $a(k)\neq 0$ for $k\in\mathbb{R}^{\ast}$ and $a(k)$ has finitely many zeros in $\mathbb{C}_{+}$. Assume that $a(k)$ has $n$ zeros and let $k_{j}$ be its zero of multiplicity $\nu_{j}$ in $\mathbb{C}_{+}$ for $j=1,\ldots,n$. Let the curve $\Gamma=\{\gamma(\xi)\in\mathbb{C}_{+}\mid\xi\in\mathbb{R}\}$ contain no pole of $u(x)$, and let $U$ be a neighborhood of $\Gamma$ in $\mathbb{C}_{+}$ containing no pole of $u(x)$, as in the proof of Proposition 2.1.
Let
$M(x;k)=\phi(x;k)e^{ikx},\quad N(x;k)=\psi(x;k)e^{ikx}.$ (3.1)
By (1.4) we have
[…]
Since $\phi(x;k)$ and $\psi(x;-k)$ are analytic in $x\in U$ and $k\in\mathbb{C}_{+}$, so are $M(x;k)$ and $N(x;-k)$. Using (1.7), we have
$M(x;k)=a(k)N(x;-k)e^{2ikx}+b(k)N(x;k).$ (3.3)
Let
$N_{j}^{r}(x):=\frac{\partial^{r}N}{\partial k^{r}}(x;k_{j}),\quad r=0,\ldots,\nu_{j}-1,\quad j=1,\ldots,n.$
Differentiating (3.3) with respect $k$ and substituting $k=k_{j}$, we obtain
$\frac{\partial^{r}M}{\partial k^{r}}(x;k_{j})=\frac{\partial^{r}}{\partial k^{r}}\bigl{(}b(k)N(x;k)\bigr{)}\Big{|}_{k=k_{j}},\quad r=0,\ldots,\nu_{j}-1,$ (3.4)
for $j=1,\ldots,n$ since $a(k)$ has a zero of multiplicity $\nu_{j}$ at $k=k_{j}$. In particular, the right hand sides of (3.4) can be represented by linear combinations of $N_{j}^{r}(x)$, $r=0,\ldots,\nu_{j}-1$. Actually, when $\nu_{j}>1$, we have
$M(x;k_{j})=b(k_{j})N_{j}^{0}(x),\quad\frac{\partial M}{\partial k}(x;k_{j})=b(k_{j})N_{j}^{1}(x)+b_{k}(k_{j})N_{j}^{0}(x),\quad\ldots.$
Define the projection operators
$\mathscr{P}^{\pm}(f)=\frac{1}{2\pi i}\int_{-\infty}^{\infty}\frac{f(\kappa)}{\kappa-(k\pm i0)}\mathrm{d}\kappa$
for $f\in L^{1}(\mathbb{R})$, where $k\in\mathbb{C}_{+}$ or $\mathbb{C}_{-}$ depending on whether the upper or lower signs are taken simultaneously. If $f_{+}$ (resp. $f_{-}$) is analytic in $\mathbb{C}_{+}$ (resp. in $\mathbb{C}_{-}$) and $f_{\pm}(k)\to 0$ as $|k|\to\infty$, then
$\mathscr{P}^{\pm}(f_{\pm})=\pm f_{\pm},\quad\mathscr{P}^{\pm}(f_{\mp})=0$
(see, e.g., Chapter 7 of [2]). Dividing both sides of (3.3) by $a(k)$ and noting (3.2), we apply $\mathscr{P}^{-}$ to the resulting equation to obtain
$N(x;k)e^{-2ikx}=1-\frac{1}{2\pi i}\sum_{j=1}^{n}\int_{|\kappa-k_{j}|=\delta}\frac{M(x;\kappa)}{(k+\kappa)a(\kappa)}\mathrm{d}\kappa$ (3.5)
since $M(x;k)$ and $N(x;-k)$ are analytic for $k\in\mathbb{C}_{+}$ and $\mathbb{C}_{-}$, respectively, and $b(k)=0$ for $k\in\mathbb{R}^{\ast}$. Using the method of residues, we see that the above integrals are represented by linear combinations of $N_{j}^{r}(x)$, $r=0,\ldots,\nu_{j}-1$. For example, when $\nu_{j}=1$,
$\frac{1}{2\pi i}\int_{|\kappa-k_{j}|=\delta}\frac{M(x;\kappa)}{(k+\kappa)a(\kappa)}\mathrm{d}\kappa=\frac{b(k_{j})}{(k+k_{j})a_{k}(k_{j})}N_{j}^{0}(x),$
and when $\nu_{j}=2$,
$\frac{1}{2\pi i}\int_{|\kappa-k_{j}|=\delta}\frac{M(x;\kappa)}{(k+\kappa)a(\kappa)}\mathrm{d}\kappa$
$=\frac{2}{(k+k_{j})a_{kk}(k_{j})}\biggl{(}b(k_{j})N_{j}^{1}(x)+\biggl{(}b_{k}(k_{j})-\frac{b(k_{j})}{k+k_{j}}-\frac{a_{kkk}(k_{j})b(k_{j})}{3a_{kk}(k_{j})}\biggr{)}N_{j}^{0}(x)\biggr{)}.$
Differentiating (3.5) with respect to $k$ up to $\nu_{j}-1$ times and setting $k=k_{j}$, we obtain a system of linear equations about $N_{j}^{r}(x)$, $r=0,\ldots,\nu_{j}-1$, $j=1,\ldots,n$, and solve it to obtain them as rational functions of $x$ and $\{e^{2ik_{j}x}\}_{j=1}^{n}$, using basic arithmetic operations. Note that
$\frac{\partial}{\partial k}(N(x;k)e^{-2ikx})=\left(\frac{\partial N}{\partial k}(x;k)-2ixN(x;k)\right)e^{-2ikx}$
and so on. Hence, we see by (3.1) and (3.5) that the Jost solution $\psi(x;k)$ is also represented by a rational function of $x$ and $\{e^{ik_{j}x}\}_{j=1}^{n}$. This means the desired result. ∎
Remark 3.1.
We show that
$u(x)=\frac{1}{\pi}\frac{\partial}{\partial x}\biggl{(}\sum_{j=1}^{n}\int_{|\kappa-k_{j}|=\delta}\frac{M(x;\kappa)}{a(\kappa)}\mathrm{d}\kappa\biggr{)}$ (3.6)
((((see, e.g., Section $9.3$ of [1]), so that $u(x)$ is also represented by rational functions of $x$ and $\{e^{ik_{j}x}\}_{j=1}^{n}$. The sign of the corresponding formula (4.12) in [26] contains an error. Note that for example, when $\nu_{j}=1$,
$\frac{1}{\pi}\int_{|\kappa-k_{j}|}\frac{M(x;\kappa)}{a(\kappa)}\mathrm{d}\kappa=\frac{2ib(k_{j})}{a_{k}(k_{j})}N_{j}^{0}(x),$
and when $\nu_{j}=2$,
$\frac{1}{\pi}\int_{|\kappa-k_{j}|}\frac{M(x;\kappa)}{a(\kappa)}\mathrm{d}\kappa$
$=\frac{4i}{a_{kk}(k_{j})}\left(b(k_{j})N_{j}^{1}(x)+\left(b_{k}(k_{j})-\frac{a_{kkk}(k_{j})b(k_{j})}{3a_{kk}(k_{j})}\right)N_{j}^{0}(x)\right).$
Since $\mathrm{Im}\,k_{j}>0$, $j=1,\ldots,n$, it follows from (3.2) and (3.6) that
$\lim_{x\to\pm\infty}u(x)=0.$
Figure 1. Potential $u(x)$ in Example 3.2.
Example 3.2.
Let $n=1$ and $\nu_{1}=2$ in the proof of Theorem 1.3. For simplicity we assume that $k_{1}=i$, $a_{kkk}(k_{1}),b_{k}(k_{1})=0$ and
$a_{kk}(k_{1})=-\tfrac{1}{2},\quad b(k_{1})=1.$ (3.7)
The system of linear equations about $N_{1}^{0}(x),N_{1}^{1}(x)$ becomes
$N_{1}^{0}(x)e^{2x}=1-2iN_{1}^{1}(x)+N_{1}^{0}(x),$
$(N_{1}^{1}(x)-2ixN_{1}^{0}(x))e^{2x}=N_{1}^{1}(x)+iN_{1}^{0}(x).$
Solving the above equation, we obtain
$N_{1}^{0}(x)=\frac{e^{2x}-1}{e^{4x}-2(2x+1)e^{2x}-1},\quad N_{1}^{1}(x)=\frac{i(2xe^{2x}+1)}{e^{4x}-2(2x+1)e^{2x}-1}.$
From (3.1) and (3.5) we have
$\psi(x;k)=N(x;k)e^{-ikx}=\biggl{(}1+\frac{4i[(2(k+i)x+i)e^{2x}+k]}{(k+i)^{2}(e^{4x}-2(2x+1)e^{2x}-1)}\biggr{)}e^{ikx}.$ (3.8)
Moreover, it follows from (3.6) that
$u(x)=-\frac{16e^{2x}(e^{2x}-1)((2x-1)e^{2x}+2x+3)}{(e^{4x}-2(2x+1)e^{2x}-1)^{2}}.$ (3.9)
See Fig. 1 for the shape of $u(x)$. It has a pole of order two at $x=0.864558\ldots$. The potential (3.9) is of a negaton-type [16, 17, 18, 24].
Let $k\in\mathbb{R}^{\ast}$. We see that
$\psi(x;k)\sim\left(\frac{k-i}{k+i}\right)^{2}e^{ikx}$
as $x\to-\infty$. Hence, it follows from (1.4) that
$\phi(x;k)=\left(\frac{k-i}{k+i}\right)^{2}\psi(x;-k),$
which yields
$a(k)=\left(\frac{k-i}{k+i}\right)^{2}$
by (1.7). Thus, $a_{k}(i)=0$ and $a_{kk}(i)=-\tfrac{1}{2}$. On the other hand, using (3.8), we see that $\psi(x;i)\sim e^{x}$ as $x\to-\infty$, and consequently $b(i)=1$. These agree with our assumption (3.7).
4. Proof of Theorem 1.4
In this section we give a proof of Theorem 1.4. Henceforth we assume that the hypotheses of Theorem 1.4 hold as well as $k\in\mathbb{R}^{\ast}$.
Letting $y=1/x$, we rewrite (1.5) as
$\frac{\mathrm{d}^{2}v}{\mathrm{d}y^{2}}+\frac{2}{y}\frac{\mathrm{d}v}{\mathrm{d}y}+\frac{k^{2}-\tilde{u}(y)}{y^{4}}v=0,$ (4.1)
where $\tilde{u}(y)=u(1/y)$. Since $(k^{2}-\tilde{u}(y))/y^{4}$ has a pole of fourth order at $y=0$, we see by a standard result on higher-order scalar linear differential equations (see, e.g., Theorem 7 of [6]) that Eq. (1.5) has an irregular singularity at infinity $x=\infty$. To obtain a necessary condition for its integrabilty, we compute its formal monodromy, exponential torus and Stokes matrices [20, 22, 25] around $y=0$ for (4.1). See these references for necessary information in the following computations.
Letting $\eta=(\mathrm{d}v/\mathrm{d}y)/v$ in (4.1), we have a Riccati equation
$\frac{\mathrm{d}\eta}{\mathrm{d}y}+\eta^{2}+\frac{2}{y}\eta+\frac{k^{2}-\tilde{u}(y)}{y^{4}}=0$ (4.2)
To obtain a formal solution to (4.2), we write
$\eta=\sum_{j=0}^{\infty}\eta_{j}y^{j-2}.$
Substituting the above expression into (4.2), we have
$\eta_{0}^{2}+k^{2}=0,\quad 2\eta_{0}\eta_{1}=0,\quad 2\eta_{0}\eta_{j}+r_{j}(\eta_{0},\ldots,\eta_{j-1})=0,\quad j\geq 2,$
where $r_{j}(\eta_{0},\ldots,\eta_{j-1})$ is a polynomial of $\eta_{0},\ldots,\eta_{j-1}$ for $j\geq 2$. Thus, we obtain formal solutions to (4.2) as
$\eta(y)=\pm\frac{ik}{y^{2}}+\cdots$
and to (4.1) as
$v(y)=\exp\left(\mp\frac{ik}{y}\right)v_{\pm}(y),$ (4.3)
where $v_{\pm}(y)$ are formal power series in $y$, and the upper or lower signs are taken simultaneously. Hence, we express have a formal fundamental matrix of (4.1) as
$V(y)=\tilde{V}(y)e^{Q(y)},$ (4.4)
where $\tilde{V}(y)$ is a $2\times 2$ formal meromorphic invertible matrix of power series in $y$ and
$Q(y)=\begin{pmatrix}-ik/y&0\\ 0&ik/y\end{pmatrix}.$
Figure 2. Singular directions $d_{\pm}$ and sectors $\Sigma_{\pm}$: The boundaries of $\Sigma_{+}$ and $\Sigma_{-}$ are plotted as solid and dotted lines..
Substitution of $re^{i\theta}$ into (4.4) yields
$V(re^{i\theta})=\tilde{V}(re^{i\theta})e^{Q(re^{i\theta})},$
so that the formal monodromy becomes
$M=V(r)^{-1}V(re^{2\pi i})=\mathrm{id}$
by $V(re^{2\pi i})=V(r)M$. The exponential torus is given by
$\mathscr{T}=\mathrm{Gal}(\mathbb{C}(e^{\mp ik/y})/\mathbb{C}(y))=\left\{\begin{pmatrix}c&0\\ 0&c^{-1}\end{pmatrix}\bigg{|}c\in\mathbb{C}^{\ast}\right\}.$
Here $\mathrm{Gal}(\mathbb{L}/\mathbb{K})$ denotes the differential Galois group, where $\mathbb{K}$ and $\mathbb{L}$ are differential fields such that $\mathbb{L}$ is a differential field extension of $\mathbb{K}$. The Stokes directions, on which $\mathrm{Re}\,(\mp ik/y)=0$, are given by $\arg y=0,\pi$. We have $\mathrm{Re}\,(-ik/y)<0$ and $\mathrm{Re}\,(ik/y)<0$ for $\arg y\in(0,\pi)$ and $(\pi,2\pi)$, respectively. So the singular directions $d_{\mp}$, which are bisectors of $(\tfrac{1}{2}\pi\mp\tfrac{1}{2}\pi,\tfrac{3}{2}\pi\mp\tfrac{1}{2}\pi)$, become $\arg y=\pi\mp\tfrac{1}{2}\pi$. See Fig. 2.
Let $a(k)\neq 0$. We take $\phi(1/y;k)$ and $\psi(1/y;-k)$ (resp. $\phi(1/y;-k)$ and $\psi(1/y;k)$) as solutions corresponding to the formal solution (4.3) with the upper (resp. lower) sign, in the sectors $\Sigma_{-}$ and $\Sigma_{+}$ containing $\mathrm{Re}\,y<0$ and $\mathrm{Re}\,y>0$, respectively, as well as the singular directions $d_{\pm}$. Near the singular direction $d_{-}$, we have
$\begin{pmatrix}\phi(1/y;k)&\phi(1/y;-k)\\ \phi^{\prime}(1/y;k)&\phi^{\prime}(1/y;-k)\end{pmatrix}=\begin{pmatrix}\psi(1/y;-k)&\psi(1/y;k)\\ \psi^{\prime}(1/y;-k)&\psi^{\prime}(1/y;k)\end{pmatrix}\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}$ (4.5)
by (1.7) and (2.3), so that the associated Stokes matrix is given by
$S_{-}=\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}.$
Near the singular direction $d_{+}$, we use (2.4) and (4.5) to obtain
$\begin{pmatrix}\psi(1/y;-k)&\psi(1/y;k)\\ \psi^{\prime}(1/y;-k)&\psi^{\prime}(1/y;k)\end{pmatrix}=\begin{pmatrix}\phi(1/y;-k)&\phi(1/y;k)\\ \phi^{\prime}(1/y;-k)&\phi^{\prime}(1/y;k)\end{pmatrix}\begin{pmatrix}a(-k)&-b(-k)\\ -b(k)&a(k)\end{pmatrix},$
so that the associated Stokes matrix is given by
$S_{+}=\begin{pmatrix}a(-k)&-b(-k)\\ -b(k)&a(k)\end{pmatrix}=S_{-}^{-1}.$
Ramis’ theorem [15, 22, 23] implies that the differential Galois group $G$ of (1.5) contains the Zariski closure of a group generated by the formal monodromy $M=\mathrm{id}$, the exponential torus $\mathscr{T}$ and the Stokes matrices $S_{\pm}$. In particular, $G$ is not finite, i.e., cases (i) and (ii) of Proposition 2.3 do not occur. Moreover, case (iv) of Proposition 2.3 does not occur since $a(k)\neq 0$. Hence, if Eq. (1.5) is integrable in the sense of differential Galois theory, then we have $b(k)$ or $b(-k)=0$ since either case (iii) or (v) of Proposition 2.3 occurs. So we have the following.
Proposition 4.1.
Let $k\in\mathbb{R}^{\ast}$ and assume that $a(k)\neq 0$. If Eq. (1.5) is solvable by quadrature, then $b(k)$ or $b(-k)=0$.
Proof of Theorem 1.4.
Suppose that Eq. (1.5) is solvable by quadrature. Since by Proposition 2.1(iii) $a(k)=0$ only at discrete points, it follows from Proposition 4.1 that $b(k)$ or $b(-k)=0$ except at the discrete points. Hence, by the identity theorem (e.g., Theorem 3.2.6 of [2]), we have $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$ since by Proposition 2.1(i) $b(k)$ is analytic in a neighborhood of $\mathbb{R}^{\ast}$ in $\mathbb{C}^{\ast}$. Thus, we complete the proof. ∎
5. Proof of Theorem 1.6
In this section we give a proof of Theorem 1.6. The proof is mainly relied to the result of Kovacic [13], which is briefly described in Appendix A.
Proof of Theorem 1.6.
Henceforth we assume that $m_{2}-m_{1}=1$. Obviously, case (c) of Proposition A.2 does not occur for $k\in\mathbb{R}^{\ast}$. So we only have to show that cases (a) and (b) of Proposition A.1 do not occur for some $k\in\mathbb{R}^{\ast}$. We rewrite (1.5) in the form (A.1) with
$r(x)=-k^{2}-u(x)$ (5.1)
and expand $u(x)$ in the Laurent series at $x=\infty$ as
$u(x)=\frac{\bar{u}_{1}}{x}+\cdots,$
where $\bar{u}_{1}\neq 0$ is a constant. Let $\mathscr{S}$ denote the set of poles of $u(x)$.
We assume that case (a) of Proposition A.1 occurs as well as $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A2, we have $d_{c}\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ in (A.2) and
$\kappa_{\infty}^{\pm}=\pm\frac{i\bar{u}_{1}}{2k}$
since $\alpha=ik$ and $\beta=-\bar{u}_{1}$. We easily see that $d_{c}\not\in\mathbb{N}_{0}$ when $k$ changes slightly even if $d_{c}\in\mathbb{N}_{0}$for some $k\in\mathbb{R}^{\ast}$ This implies that case (a) of Proposition A.1 does not occur in an open interval of $k\in\mathbb{R}^{\ast}$ since $\mathscr{S}$ is finite so that the number of a finite sequence $c$ of the symbols $\pm$ with $d_{c}\in\mathbb{N}_{0}$ is finite.
We next assume that case (b) of Proposition A.1 occurs for an open interval of $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A3, we have $d_{e}\in\mathbb{N}_{0}$ in (A.3) and there exists a monic polynomial $P(x)$ of order $d_{e}$ satisfying (A.5). Since
$\lim_{x\to\pm\infty}(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))=4k^{2},$
we easily see that $d_{e}\neq 0$. Let $d=d_{e}\geq 1$. Since Eq. (A.5) holds, we see that
$\begin{split}&(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x)),\\ &(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))x\\ &\qquad+(3\theta(x)^{2}+3\theta^{\prime}(x)+4(k^{2}+u(x))),\\ &(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))x^{2}\\ &\quad+2(3\theta(x)^{2}+3\theta^{\prime}(x)+4(k^{2}+u(x)))x+6\theta(x),\\ &(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))x^{j}\\ &\quad+j(3\theta(x)^{2}+3\theta^{\prime}(x)+4(k^{2}+u(x)))x^{j-1}+3j(j-1)\theta(x)x^{j-2}\\ &\quad+j(j-1)(j-2)x^{j-3},\quad j=3,\ldots,d,\end{split}$ (5.2)
are linearly dependent, on the interval of $k$, where the last equation is eliminated if $d=2$, and the last two equations are eliminated if $d=1$. Hence,
$\theta(x),\quad(\theta(x)x+j)x^{j-1},\quad j=1,\ldots,d,$
are linearly dependent. If $d=1$, then
$(\theta(x)x+1)+q_{0}\theta(x)=\theta(x)(x+q_{0})+1=0.$
If $d\geq 2$, then
$\theta(x),\quad(\theta(x)x+j)x^{j-1},\quad j=1,\ldots,d-1,$
are linearly independent, so that for certain constants $q_{j}$, $j=0,\ldots,d-1$,
$(\theta(x)x+d)x^{d-1}+\sum_{j=1}^{d-1}q_{j}(\theta(x)x+j)x^{j-1}+q_{0}\theta(x)$
$=\theta(x)\sum_{j=0}^{d}q_{j}x^{j}+\sum_{j=1}^{d}jq_{j}x^{j-1}=0,$
where $q_{d}=1$. Thus, we have
$\theta(x)=-\frac{q^{\prime}(x)}{q(x)},\quad q(x)=\sum_{j=0}^{d}q_{j}x^{j},$ (5.3)
and
$P(x)=\sum_{j=0}^{d}q_{j}x^{j}=q(x),$ (5.4)
since the functions given by (5.2) are linearly dependent on the interval of $k$. Substituting (5.3) and (5.4) into (A.5), we obtain
$q(x)u^{\prime}(x)+4q^{\prime}(x)u(x)=0,$
which yields
$u(x)=Cq(x)^{-4},$
where $C\neq 0$ is a constant. This means that $m_{2}-m_{1}=-4$ and yields a contradiction. Thus we complete the proof. ∎
Remark 5.1.
In the above proof, an essential role was played by the fact that poles of $r(x)$ given by (5.1) do not depend on $k$.
Example 5.2.
Consider the rational potential
$u(x)=-\frac{5}{16x^{2}}-\frac{5}{16(x-1)^{2}}-\frac{7}{8x}+\frac{5}{24(x-1)}$
and let $r(x)=-k^{2}-u(x)$. Equation (1.5) satisfies $m_{2}-m_{1}=1$ and only has singularities of order two at $x=0$ and $1$. Following the recipe of Appendix A.2, we have $\beta=5/16$ and
$\kappa_{s}^{\pm}=\tfrac{1}{2}\pm\tfrac{3}{4}$
for $s=0,1$, and
$\kappa_{\infty}^{\pm}=\mp\frac{19i}{48k}$
since
$\bar{u}_{1}=\lim_{x\to\infty}xu(x)=-\tfrac{19}{24}.$
Hence, case (a) of Proposition A.1 does not occur for any $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A.3, we have $\beta=5/16$ and
$E_{s}=\{-1,2,5\}$
for $s=0,1$. We see that $d_{e}=1$ for $e_{0},e_{1}=-1$ is only a non-negative integer, so that
$\theta(x)=-\frac{1}{2x}-\frac{1}{2(x-1)}.$
Substituting $P(x)=x+p_{0}$ into (A.5), we obtain
$\frac{1}{x^{2}(x-1)^{3}}\bigl{(}-(k^{2}p_{0}+2k^{2}-\tfrac{4}{3})x^{4}+(2(5k^{2}+2)p_{0}+4k^{2}-1)x^{3}$
$-\tfrac{1}{3}(8(3k^{2}-5)p_{0}+2(3k^{2}+5))x^{2}+\tfrac{1}{3}(2(3k^{2}+20)p_{0}+4)x+4p_{0}+1\bigr{)}=0.$
Hence, case (b) of Proposition A.1 holds with $p_{0}=-\tfrac{1}{4}$, so that Eq. (1.5) is solvable by quadrature, if and only if $k=\pm\tfrac{2}{3}\sqrt{3}$.
Let $k=\pm\tfrac{2}{3}\sqrt{3}$. We have
$\hat{\theta}(x)=-\frac{1}{2x}-\frac{1}{2(x-1)}+\frac{4}{4x-1},$
so that the integrand in (A.6) satisfies
$\tfrac{1}{2}\left(-\hat{\theta}(x)\pm\sqrt{-4(k^{2}+u(x))-\hat{\theta}(x)^{2}-2\hat{\theta}^{\prime}(x)}\right)\to\pm ik\left(1-\frac{19}{48k^{2}x}\right)+O(x^{-2})$
as $|x|\to\infty$ on $\mathbb{R}$. This implies that Eq. (1.5) has two linearly independent solutions such that
$v(x)\to x^{-19i/48k}e^{ikx}\mbox{ and }x^{19i/48k}e^{-ikx}\quad\mbox{as $x\to\pm\infty$}.$
Thus, Eq. (1.5) does not have the Jost solutions satisfying (1.4) even when it is solvable by quadrature.
Acknowledgements
This work was partially supported by the JSPS KAKENHI Grant Number JP22H01138.
Data Availability
Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Conflict of Interest Statement
The author has no conflict of interest to disclose.
Appendix A Kovacic’s Result
In this appendix we review necessary parts of Kovacic [13] on second-order linear differential equations on $\mathbb{C}(x)$.
A.1. Fundamental results
Consider second-order linear differential equations of the form
$v_{xx}=r(x)v,$ (A.1)
where $r(x)\in\mathbb{C}(x)\setminus\mathbb{C}$. Let $G$ be the differential Galois group of (A.1) over $\mathbb{C}(x)$. We easily see that $G$ is an algebraic subgroup of $\mathrm{SL}(2,\mathbb{C})$ (see, e.g., Chapter 6 of [12]). We have the following four cases for (A.1) (see [10, 13] for a proof).
Proposition A.1 (Kaplansky [12], Kovacic[13]).
One of the following four cases occurs::\,::
  1. (a)
    $G$ is triangularizable;absent\,;;
  2. (b)
    $G$ is conjugate to a subgroup of
    $D^{\dagger}=\left\{\begin{pmatrix}c&0\\ 0&c^{-1}\end{pmatrix}\bigg{|}\,c\in\mathbb{C}^{\ast}\right\}\cup\left\{\begin{pmatrix}0&c\\ -c^{-1}&0\end{pmatrix}\bigg{|}\,c\in\mathbb{C}^{\ast}\right\},$
    but it is not triangularizable;absent\,;;
  3. (c)
    $G$ is finite, but neither triangulariable nor conjugate to $D^{\dagger};$
  4. (d)
    $G=\mathrm{SL}(2,\mathbb{C})$.
If one of cases (a)-(c) of Proposition A.1 occurs, then the connected identity component $G^{0}$ of $G$ is solvable and Eq. (A.1) is solvable. On the other hand, if cases (a)-(c) do not occur, then $G=G^{0}=\mathrm{SL}(2,\mathbb{C})$, i.e., Eq. (A.1) is not integrable in the sence of differential Galois theory.
Let $r(x)=r_{1}(x)/r_{2}(x)$, where $r_{1}(x),r_{2}(x)\in\mathbb{C}[z]$ are relatively prime. Poles of $r(x)$ are zeros of $r_{2}(x)$ and their orders are the multiplicities of the zeros of $r_{2}(x)$. Let $\mathscr{S}\subset\mathbb{C}$ denote the set of all poles of $r(x)$. By the order of $r(x)$ at $\infty$, we also mean the order of $\infty$ as a zero of $r(x)$, i.e., $\deg r_{2}(x)-\deg r_{1}(x)$. We have the following necessary conditions for cases (a)-(c) in Proposition A.1 to occur (see [10, 13] for a proof).
Proposition A.2 (Kovacic [13]).
The following conditions are necessary for the corresponding cases in Proposition A.1 to occur::\,::
  • (a)
    No pole of $r(x)$ is of odd order greater than one, and the order of $r(x)$ at $\infty$ is not an odd less than two;absent\,;;
  • (b)
    There exists a pole of $r(x)$ such that its order is two or an odd greater than two;absent\,;;
  • (c)
    No pole of $r(x)$ is of order greater than two and the order of $r(x)$ at $\infty$ is greater than one.
Note that cases (a), (b) and (c) of Proposition A.1 may not occur even though conditions (a), (b) and (c) hold, respectively. In particular, Proposition A.2 only states fewer conditions for case (c) than the original version in [13] since they are enough for our application in Section 5. Kovacic [13] also gave an algorithm for finding a “closed-form” solution and detecting whether cases (a)-(c) in Proposition A.1 really occur or not under conditions (a), (b) and (c) of Proposition A.1, respectively. In the following we describe necessary parts of his algorithm for cases (a) and (b) in the setting required in Section 5: The order of $r(x)$ at $\infty$ is zero. See [10, 13] for the full version of his algorithm and its proof.
A.2. Case (a)
We assume that condition (a) in Proposition A.2 holds as well as the order of $r(x)$ at $\infty$ is zero. When $g(x)$ has a pole of order $\ell\geq 2$ at $x=s$ and is expanded in a Laurent series
$g(x)=\sum_{j\geq-\ell}g_{j}(x-s)^{j}$
there, we write
$[g(x)]_{s}=\sum_{-\ell\leq j\leq-2}g_{j}(x-s)^{j}.$
Similarly, when $g(x)$ is of order $\ell$ at $\infty$ and is expanded in a Laurent series
$g(x)=\sum_{j\leq\ell}g_{j}x^{j}$
there, we write
$[g(x)]_{\infty}=\sum_{0\leq j\leq\ell}g_{j}x^{j}.$
For $s\in\mathscr{S}\cup\{\infty\}$ we define a rational function $\omega_{s}(x)$ and two constants $\kappa_{s}^{\pm}$ as follows:
  1. ($s_{1}$)
    If $s\in\mathscr{S}$ is of order $1$, then $\kappa_{s}^{\pm}=1$;
  2. ($s_{2}$)
    If $s\in\mathscr{S}$ is of order $2$, then
    $\kappa_{s}^{\pm}=\tfrac{1}{2}\pm\tfrac{1}{2}\sqrt{1+4\beta},$
    where $\beta$ is the coefficient of $1/(x-s)^{2}$ in the partial fraction expansion of $r(x);$
  3. ($s_{3}$)
    If $s\in\mathscr{S}$ is of order $22\ell>$ 2, then
    $\kappa_{s}^{\pm}=\tfrac{1}{2}\left(\pm\frac{\beta}{\alpha}+\ell\right),$
    where $\alpha$ and $\beta$ are the coefficients of $(x-s)^{-\ell}$ and $(x-s)^{-(\ell+1)}$ in the Laurent series expansions of $\sqrt{r(x)}$ and $r(x)-([\sqrt{r(x)}]_{s})^{2}$, respectively;
  1. ($\infty_{1}$)
    $\omega_{\infty}(x)=[\sqrt{r(x)}]_{\infty}$ and
    $\kappa_{\infty}^{\pm}=\pm\frac{\beta}{2\alpha},$
    where $\alpha$ and $\beta$ are the coefficients of $x^{0}$ and $x^{-1}$ in the Laurent series expansions of $\sqrt{r(x)}$ and $r(x)$, respectively.
For each sequence $c=\{c(s)\}_{s\in\mathscr{S}\cup\{\infty\}}$ with $c(s)=+$ or --, define
$d_{c}=\kappa_{\infty}^{c(\infty)}-\sum_{s\in\mathscr{S}}\kappa_{s}^{c(s)}.$ (A.2)
If $d_{c}$ is not a non-negative integer for any sequence $c$, then case (a) of Proposition A.1 does not occur. We also have a further recipe to determine whether it occurs or not when $d_{c}$ is a non-negative integer, although it is not given here since it is not necessary for our purpose. The reader who is interested in it should consult [10, 13].
A.3. Case (b)
We assume that condition (b) in Proposition A.2 holds as well as the order of $r(x)$ at $\infty$ is zero. For $s\in\mathscr{S}\cup\{\infty\}$ we define a set $E_{s}\subset\mathbb{Z}$ as follows:
  1. ($s_{1}$)
    If $s\in\mathscr{S}$ is of order $1$, then $E_{s}=\{4\}$;
  2. ($s_{2}$)
    If $s\in\mathscr{S}$ is of order $2$, then
    $E_{s}=\{2+l\sqrt{1+4\beta}\mid l=0,\pm 2\}\cap\mathbb{Z},$
    where $\beta$ is the coefficient of $1/(x-s)^{2}$ in the partial fraction expansion of $r\,;$
  3. ($s_{3}$)
    If $s\in\mathscr{S}$ is of order $j>2$, then $E_{s}=\{j\}$.
For each tuple $e=\{e_{s}\in E_{s}\mid s\in\mathscr{S}\}$, let
$d_{e}=-\tfrac{1}{2}\sum_{s\in\mathscr{S}}e_{s}.$ (A.3)
If $d_{e}$ is not a non-negative integer for any $e$, then case (b) of Proposition A.1 does not occur.
Let $E=\{e\mid d_{e}\in\mathbb{N}_{0}\}$ and let
$\theta(x)=\tfrac{1}{2}\sum_{s\in\mathscr{S}}\frac{e_{s}}{x-s}$ (A.4)
for $e\in E$. By (A.3) and (A.4) we see that the Laurent series of $\theta(x)$ at $\infty$ is given by
$\theta(x)=-\frac{d_{e}}{x}+\cdots.$
If there exists a monic polynomial $P(x)$ of degree $d_{e}$ satisfying
$P^{\prime\prime\prime}+3\theta(x)P^{\prime\prime}+(3\theta(x)^{2}+3\theta^{\prime}(x)-4r(x))P^{\prime}$
$+(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}-4r(x)\theta(x)-2r^{\prime}(x))P=0,$ (A.5)
then case (b) of Proposition A.1 occurs and
$v(x)=\exp\left(\int\tfrac{1}{2}\left(-\hat{\theta}(x)\pm\sqrt{4r(x)-\hat{\theta}(x)^{2}-2\hat{\theta}^{\prime}(x)}\right)\mathrm{d}x\right)$ (A.6)
are solutions to (A.1), where $\hat{\theta}(x)=\theta(x)+P^{\prime}(x)/P(x)$. If such a polynomial is not found for any sequence $e$, then it does not occur.

### v2

@@
 Remark 1.5.
+  • (i)
     The scattering coefficient $a(k)$ may have no zero in $\mathbb{C}_{+}$ when condition (A1) holds and $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$.
@@
 In such a case, Theorem 1.3 does not apply.
+  • (ii)
+    Statements similar to those of Theorems 1.3 and 1.4 are proven for the ZS systems in [30].
 Furthermore, we consider the case in which $u(x)$ is a rational function and satisfies (1.9) but does not (1.8).
@@
 We give proofs of Theorems 1.3, 1.4 and 1.6 in Sections 3, 4 and 5, respectively.
-The result of Kovacic [13], which plays a crucial role in Section 5, is briefly reviewed in Appendix A.
+Fundamental information on local differential Galois groups near irregular singularities, which is needed in Section 4, is provided in Appendix A, and the result of Kovacic [13], which plays a crucial role in Section 5, is briefly reviewed in Appendix B.
 2. Preliminary Results
@@
 Remark 2.2.
-The conclusion in Proposition 2.1(iv) holds under a weaker condition. For example, if $\lim_{k\to 0}b(k)b(-k)\to b_{0}$ $(\neq-1)$, then it holds.
+The conclusion in Proposition 2.1(iv) holds under a weaker condition. For example, if $\lim_{k\to 0}b(k)b(-k)=b_{0}$ $(\neq-1)$, then it holds.
 Let $G$ be the differential Galois group of (1.5).
@@
     $\mathscr{G}=\mathscr{G}^{0}=\mathrm{SL}(2,\,\mathbb{C})$.
-The system (1.5) is integrable in the sense of differential Galois theory unless $G$ is of type (vi). Proposition 2.3 plays a key role in the proof of Theorem 1.4 in Section 4.
+The system (1.5) is solvable by quadrature unless $G$ is of type (vi). Proposition 2.3 plays a key role in the proof of Theorem 1.4 in Section 4.
 3. Proof of Theorem 1.3
@@
 (1.5) has an irregular singularity at infinity $x=\infty$.
-To obtain a necessary condition for its integrabilty, we compute its formal monodromy, exponential torus and Stokes matrices [20, 22, 25] around $y=0$ for (4.1).
-See these references for necessary information in the following computations.
+To obtain a necessary condition for its integrabilty, we compute its formal monodromy, exponential torus and Stokes matrices around $y=0$ for (4.1).
+See Appendix A for necessary information, and see also [21, 23, 28] for more details.
 Letting $\eta=(\mathrm{d}v/\mathrm{d}y)/v$ in (4.1), we have a Riccati equation
-$\frac{\mathrm{d}\eta}{\mathrm{d}y}+\eta^{2}+\frac{2}{y}\eta+\frac{k^{2}-\tilde{u}(y)}{y^{4}}=0$ (4.2)
+$\frac{\mathrm{d}\eta}{\mathrm{d}y}+\eta^{2}+\frac{2}{y}\eta+\frac{k^{2}-\tilde{u}(y)}{y^{4}}=0.$ (4.2)
 To obtain a formal solution to (4.2), we write
@@
 Figure 2. Singular directions $d_{\pm}$ and sectors $\Sigma_{\pm}$: The boundaries of $\Sigma_{+}$ and $\Sigma_{-}$ are plotted as solid and dotted lines..
-Substitution of $re^{i\theta}$ into (4.4) yields
-$V(re^{i\theta})=\tilde{V}(re^{i\theta})e^{Q(re^{i\theta})},$
+Substitution of $ye^{2\pi i}$ into (4.4) yields
+$V(ye^{i\theta})=\tilde{V}(ye^{2\pi i})e^{Q(ye^{2\pi i})},$
 so that the formal monodromy becomes
-$M=V(r)^{-1}V(re^{2\pi i})=\mathrm{id}$
-by $V(re^{2\pi i})=V(r)M$. The exponential torus is given by
-$\mathscr{T}=\mathrm{Gal}(\mathbb{C}(e^{\mp ik/y})/\mathbb{C}(y))=\left\{\begin{pmatrix}c&0\\ 0&c^{-1}\end{pmatrix}\bigg{|}c\in\mathbb{C}^{\ast}\right\}.$
+$\hat{M}=V(y)^{-1}V(ye^{2\pi i})=\mathrm{id}$
+by $V(ye^{2\pi i})=V(y)\hat{M}$. The exponential torus is given by
+$\mathscr{T}=\mathscr{T}:=\mathrm{Gal}\left(\mathbb{C}((y))(e^{ik/y},e^{-ik/y})/\mathbb{C}((y))\right)=\left\{\begin{pmatrix}c&0\\ 0&c^{-1}\end{pmatrix}\bigg{|}c\in\mathbb{C}^{\ast}\right\}.$
 Here $\mathrm{Gal}(\mathbb{L}/\mathbb{K})$ denotes the differential Galois group, where $\mathbb{K}$ and $\mathbb{L}$ are differential fields such that $\mathbb{L}$ is a differential field extension of $\mathbb{K}$.
@@
 2.
-Let $a(k)\neq 0$.
-We take $\phi(1/y;k)$ and $\psi(1/y;-k)$ (resp.
-$\phi(1/y;-k)$ and $\psi(1/y;k)$) as solutions corresponding to the formal solution (4.3) with the upper (resp.
-lower) sign, in the sectors $\Sigma_{-}$ and $\Sigma_{+}$ containing $\mathrm{Re}\,y<0$ and $\mathrm{Re}\,y>0$, respectively, as well as the singular directions $d_{\pm}$.
-Near the singular direction $d_{-}$, we have
-$\begin{pmatrix}\phi(1/y;k)&\phi(1/y;-k)\\ \phi^{\prime}(1/y;k)&\phi^{\prime}(1/y;-k)\end{pmatrix}=\begin{pmatrix}\psi(1/y;-k)&\psi(1/y;k)\\ \psi^{\prime}(1/y;-k)&\psi^{\prime}(1/y;k)\end{pmatrix}\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}$ (4.5)
-by (1.7) and (2.3), so that the associated Stokes matrix is given by
-$S_{-}=\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}.$
-Near the singular direction $d_{+}$, we use (2.4) and (4.5) to obtain
-$\begin{pmatrix}\psi(1/y;-k)&\psi(1/y;k)\\ \psi^{\prime}(1/y;-k)&\psi^{\prime}(1/y;k)\end{pmatrix}=\begin{pmatrix}\phi(1/y;-k)&\phi(1/y;k)\\ \phi^{\prime}(1/y;-k)&\phi^{\prime}(1/y;k)\end{pmatrix}\begin{pmatrix}a(-k)&-b(-k)\\ -b(k)&a(k)\end{pmatrix},$
-so that the associated Stokes matrix is given by
-$S_{+}=\begin{pmatrix}a(-k)&-b(-k)\\ -b(k)&a(k)\end{pmatrix}=S_{-}^{-1}.$
-Ramis’ theorem [15, 22, 23] implies that the differential Galois group $G$ of (1.5) contains the Zariski closure of a group generated by the formal monodromy $M=\mathrm{id}$, the exponential torus $\mathscr{T}$ and the Stokes matrices $S_{\pm}$.
+By Proposition  A.1, we can write the Stokes matrices as
+$S_{-}=\begin{pmatrix}1&\alpha_{-}\\ 0&1\end{pmatrix}\quad\mbox{and}\quad S_{+}=\begin{pmatrix}1&0\\ \alpha_{+}&1\end{pmatrix}$ (4.5)
+for the singular directions $d_{-}$ and $d_{+}$, respectively, where $\alpha_{\pm}$ are constants which may be zero.
+Suppose that $a(k),a(-k)\neq 0$. Let
+$\Phi(y;k)=\begin{pmatrix}\phi(y^{-1};k)&\phi(y^{-1};-k)\\ -y^{-2}\phi^{\prime}(y^{-1};k)&-y^{-2}\phi^{\prime}(y^{-1};-k)\end{pmatrix},$
+$\Psi(y;k)=\begin{pmatrix}\psi(y^{-1};-k)&\psi(y^{-1};k)\\ -y^{-2}\psi^{\prime}(y^{-1};-k)&-y^{-2}\psi^{\prime}(y^{-1};k)\end{pmatrix},$
+which are fundamental matrices to (4.1). By (1.7) and (2.3),
+$\Phi(y;k)=\Psi(y;k)B(k),\quad B(k)=\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}.$ (4.6)
+Lemma 4.1.
+If $\alpha_{-}$ and $\alpha_{+}=0$, then $b(-k)$ and $b(k)=0$, respectively.
+Proof.
+We prove that if $\alpha_{-}=0$, then $b(-k)=0$. The remaining part can be proven similarly.
+First, by (1.4), we can write the formal fundamental matrix as
+$V(y)=\left(\begin{array}[]{cc}\begin{matrix}*\\ *\end{matrix}\quad\Phi(y;k)\begin{pmatrix}0\\ c_{-}\end{pmatrix}\end{array}\right)\quad\mbox{and}\quad\left(\begin{array}[]{cc}\begin{matrix}*\\ *\end{matrix}\quad\Psi(y;k)\begin{pmatrix}0\\ c_{+}\end{pmatrix}\end{array}\right)$
+in the directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$ near $d_{-}$ in $\Sigma_{-}$ and $\Sigma_{+}$, respectively, where $c_{\pm}\in\mathbb{C}$ are nonzero constants.
+If $\alpha_{-}=0$, then we use (4.5) and (4.6) to obtain
+$b(-k)c_{-}=0,\quad a(-k)c_{-}-c_{+}=0,$
+by the definition of the Stokes matrix. Thus, we obtain the desired result. ∎
+Theorem A.2 implies that the differential Galois group $G$ of (1.5) contains the Zariski closure of a group generated by the formal monodromy $\hat{M}=\mathrm{id}$, the exponential torus $\mathscr{T}$ and the Stokes matrices $S_{\pm}$.
 In particular, $G$ is not finite, i.e., cases (i) and (ii) of Proposition 2.3 do not occur.
@@
 Hence, if Eq.
-(1.5) is integrable in the sense of differential Galois theory, then we have $b(k)$ or $b(-k)=0$ since either case (iii) or (v) of Proposition 2.3 occurs.
+(1.5) is integrable in the sense of differential Galois theory, then we have have $\alpha_{-}$ or $\alpha_{+}=0$, so that by Lemma 4.1 $b(k)$ or $b(-k)=0$, since either case (iii) or (v) of Proposition 2.3 occurs.
 So we have the following.
-Proposition 4.1.
+Lemma 4.2.
 Let $k\in\mathbb{R}^{\ast}$ and assume that $a(k)\neq 0$. If Eq. (1.5) is solvable by quadrature, then $b(k)$ or $b(-k)=0$.
@@
 (1.5) is solvable by quadrature.
-Since by Proposition 2.1(iii) $a(k)=0$ only at discrete points, it follows from Proposition 4.1 that $b(k)$ or $b(-k)=0$ except at the discrete points.
+Since by Proposition 2.1(iii) $a(k)=0$ only at discrete points, it follows from Lemma 4.2 that $b(k)$ or $b(-k)=0$ except at the discrete points.
 Hence, by the identity theorem (e.g., Theorem 3.2.6 of [2]), we have $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$ since by Proposition 2.1(i) $b(k)$ is analytic in a neighborhood of $\mathbb{R}^{\ast}$ in $\mathbb{C}^{\ast}$.
@@
 5. Proof of Theorem 1.6
-In this section we give a proof of Theorem 1.6. The proof is mainly relied to the result of Kovacic [13], which is briefly described in Appendix A.
+In this section we give a proof of Theorem 1.6. The proof is mainly relied to the result of Kovacic [13], which is briefly described in Appendix B.
 Proof of Theorem 1.6.
 Henceforth we assume that $m_{2}-m_{1}=1$.
-Obviously, case (c) of Proposition A.2 does not occur for $k\in\mathbb{R}^{\ast}$.
-So we only have to show that cases (a) and (b) of Proposition A.1 do not occur for some $k\in\mathbb{R}^{\ast}$.
-We rewrite (1.5) in the form (A.1) with
+Obviously, case (c) of Proposition B.2 does not occur for $k\in\mathbb{R}^{\ast}$.
+So we only have to show that cases (a) and (b) of Proposition B.1 do not occur for some $k\in\mathbb{R}^{\ast}$.
+We rewrite (1.5) in the form (B.1) with
 $r(x)=-k^{2}-u(x)$ (5.1)
@@
 where $\bar{u}_{1}\neq 0$ is a constant. Let $\mathscr{S}$ denote the set of poles of $u(x)$.
-We assume that case (a) of Proposition A.1 occurs as well as $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A2, we have $d_{c}\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ in (A.2) and
+We assume that case (a) of Proposition B.1 occurs as well as $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix B2, we have $d_{c}\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ in (B.2) and
 $\kappa_{\infty}^{\pm}=\pm\frac{i\bar{u}_{1}}{2k}$
 since $\alpha=ik$ and $\beta=-\bar{u}_{1}$.
-We easily see that $d_{c}\not\in\mathbb{N}_{0}$ when $k$ changes slightly even if $d_{c}\in\mathbb{N}_{0}$for some $k\in\mathbb{R}^{\ast}$ This implies that case (a) of Proposition A.1 does not occur in an open interval of $k\in\mathbb{R}^{\ast}$ since $\mathscr{S}$ is finite so that the number of a finite sequence $c$ of the symbols $\pm$ with $d_{c}\in\mathbb{N}_{0}$ is finite.
-We next assume that case (b) of Proposition A.1 occurs for an open interval of $k\in\mathbb{R}^{\ast}$.
-Following the recipe of Appendix A3, we have $d_{e}\in\mathbb{N}_{0}$ in (A.3) and there exists a monic polynomial $P(x)$ of order $d_{e}$ satisfying (A.5).
+We easily see that $d_{c}\not\in\mathbb{N}_{0}$ when $k$ changes slightly even if $d_{c}\in\mathbb{N}_{0}$for some $k\in\mathbb{R}^{\ast}$ This implies that case (a) of Proposition B.1 does not occur in an open interval of $k\in\mathbb{R}^{\ast}$ since $\mathscr{S}$ is finite so that the number of a finite sequence $c$ of the symbols $\pm$ with $d_{c}\in\mathbb{N}_{0}$ is finite.
+We next assume that case (b) of Proposition B.1 occurs for an open interval of $k\in\mathbb{R}^{\ast}$.
+Following the recipe of Appendix B3, we have $d_{e}\in\mathbb{N}_{0}$ in (B.3) and there exists a monic polynomial $P(x)$ of order $d_{e}$ satisfying (B.5).
 Since
 $\lim_{x\to\pm\infty}(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))=4k^{2},$
-we easily see that $d_{e}\neq 0$. Let $d=d_{e}\geq 1$. Since Eq. (A.5) holds, we see that
+we easily see that $d_{e}\neq 0$. Let $d=d_{e}\geq 1$. Since Eq. (B.5) holds, we see that
 $\begin{split}&(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x)),\\ &(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))x\\ &\qquad+(3\theta(x)^{2}+3\theta^{\prime}(x)+4(k^{2}+u(x))),\\ &(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))x^{2}\\ &\quad+2(3\theta(x)^{2}+3\theta^{\prime}(x)+4(k^{2}+u(x)))x+6\theta(x),\\ &(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}+4(k^{2}+u(x))\theta(x)+2u^{\prime}(x))x^{j}\\ &\quad+j(3\theta(x)^{2}+3\theta^{\prime}(x)+4(k^{2}+u(x)))x^{j-1}+3j(j-1)\theta(x)x^{j-2}\\ &\quad+j(j-1)(j-2)x^{j-3},\quad j=3,\ldots,d,\end{split}$ (5.2)
@@
 $P(x)=\sum_{j=0}^{d}q_{j}x^{j}=q(x),$ (5.4)
-since the functions given by (5.2) are linearly dependent on the interval of $k$. Substituting (5.3) and (5.4) into (A.5), we obtain
+since the functions given by (5.2) are linearly dependent on the interval of $k$. Substituting (5.3) and (5.4) into (B.5), we obtain
 $q(x)u^{\prime}(x)+4q^{\prime}(x)u(x)=0,$
@@
 $u(x)=-\frac{5}{16x^{2}}-\frac{5}{16(x-1)^{2}}-\frac{7}{8x}+\frac{5}{24(x-1)}$
-and let $r(x)=-k^{2}-u(x)$. Equation (1.5) satisfies $m_{2}-m_{1}=1$ and only has singularities of order two at $x=0$ and $1$. Following the recipe of Appendix A.2, we have $\beta=5/16$ and
+and let $r(x)=-k^{2}-u(x)$. Equation (1.5) satisfies $m_{2}-m_{1}=1$ and only has singularities of order two at $x=0$ and $1$. Following the recipe of Appendix B.2, we have $\beta=5/16$ and
 $\kappa_{s}^{\pm}=\tfrac{1}{2}\pm\tfrac{3}{4}$
@@
 $\bar{u}_{1}=\lim_{x\to\infty}xu(x)=-\tfrac{19}{24}.$
-Hence, case (a) of Proposition A.1 does not occur for any $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A.3, we have $\beta=5/16$ and
+Hence, case (a) of Proposition B.1 does not occur for any $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix B.3, we have $\beta=5/16$ and
 $E_{s}=\{-1,2,5\}$
@@
 $\theta(x)=-\frac{1}{2x}-\frac{1}{2(x-1)}.$
-Substituting $P(x)=x+p_{0}$ into (A.5), we obtain
+Substituting $P(x)=x+p_{0}$ into (B.5), we obtain
 $\frac{1}{x^{2}(x-1)^{3}}\bigl{(}-(k^{2}p_{0}+2k^{2}-\tfrac{4}{3})x^{4}+(2(5k^{2}+2)p_{0}+4k^{2}-1)x^{3}$
 $-\tfrac{1}{3}(8(3k^{2}-5)p_{0}+2(3k^{2}+5))x^{2}+\tfrac{1}{3}(2(3k^{2}+20)p_{0}+4)x+4p_{0}+1\bigr{)}=0.$
-Hence, case (b) of Proposition A.1 holds with $p_{0}=-\tfrac{1}{4}$, so that Eq. (1.5) is solvable by quadrature, if and only if $k=\pm\tfrac{2}{3}\sqrt{3}$.
+Hence, case (b) of Proposition B.1 holds with $p_{0}=-\tfrac{1}{4}$, so that Eq. (1.5) is solvable by quadrature, if and only if $k=\pm\tfrac{2}{3}\sqrt{3}$.
 Let $k=\pm\tfrac{2}{3}\sqrt{3}$. We have
 $\hat{\theta}(x)=-\frac{1}{2x}-\frac{1}{2(x-1)}+\frac{4}{4x-1},$
-so that the integrand in (A.6) satisfies
+so that the integrand in (B.6) satisfies
 $\tfrac{1}{2}\left(-\hat{\theta}(x)\pm\sqrt{-4(k^{2}+u(x))-\hat{\theta}(x)^{2}-2\hat{\theta}^{\prime}(x)}\right)\to\pm ik\left(1-\frac{19}{48k^{2}x}\right)+O(x^{-2})$
@@
 The author has no conflict of interest to disclose.
-Appendix A Kovacic’s Result
+Appendix A Local Differential Galois Groups near Irregular Singularities
+In this appendix we review basic information on local differential Galois groups, which consist of formal monodromies, exponential tori and Stokes matrices, near irregularity singular points.
+See [21, 23, 28] for the details.
+Recall some terminologies of the differential Galois theory in Section 3 of [29] (see also [10, 23]) if necessary in the following.
+Consider $n$-dimensional linear systems of the form
+$\frac{\mathrm{d}\zeta}{\mathrm{d}y}=A(y)\zeta$ (A.1)
+with an irregular singularity at $y=0$ in $\mathbb{C}$, where $A(y)\in\mathrm{gl}(n,\mathbb{C}((y)))$, and $\mathbb{C}((y))$ denotes the field of the formal Laurent series over $\mathbb{C}$.
+In a standard way, Eq.
+(4.1) can be written in the form (A.1) easily.
+The linear system (A.1) has a formal fundamental matrix of the form
+$Y(y)=\hat{Y}(y)y^{L}e^{Q(1/z)}$ (A.2)
+like (4.4), where
+  • $y^{\ell}=z$ for some $\ell=1,\ldots,n$;
+  • $\hat{Y}(y)\in\mathrm{GL}(n;\mathbb{C}((y)))$;
+  • $L\in\mathrm{gl}(n;\mathbb{C})$;
+  • $Q(1/z)=\operatorname{diag}(q_{1}(z),\ldots,q_{n}(z))$, $q_{j}(z)\in\frac{1}{z}\mathbb{C}[\frac{1}{z}]$
+(see, e.g., [16]).
+So we obtain a formal Picard-Vessiot extension $\mathbb{L}\subset\mathbb{K}=\mathbb{C}((y))$, where $\mathbb{L}$ is a differential field generated by entries of $Y(y)$ and having $\mathbb{C}$ as its field of constants.
+Let $\hat{\mu}$ be a $\mathbb{K}$-automorphism of $\mathbb{L}$ satisfying
+  • $\hat{\mu}(y^{\alpha_{j}})=e^{2\pi i\alpha_{j}}y^{\alpha_{j}}$ for any $j=1,\ldots,n$;
+  • $\hat{\mu}(\log y)=\log y+2\pi i$;
+  • $\hat{\mu}(\exp(q_{j}(y^{1/\ell})))=\exp(q_{j}(e^{2\pi i/\ell}y^{1/\ell})$,
+where $\alpha_{j}$, $j=1,\ldots,n$, denote the eigenvalues of $L$.
+Since $\hat{\mu}$ commutes the differentiation, $\hat{\mu}(Y(y))$ is also a formal fundamental matrix.
+Hence, there exists a matrix $\hat{M}\in\mathrm{GL}(n;\mathbb{C})$ such that
+$\hat{\mu}(Y(y))=Y(y)\hat{M}.$
+We call $\hat{M}$ the formal monodromy of (A.1). Thus, $\hat{M}$ can be computed by substituting $ye^{i\theta}$ into $Y(y)$, as in Section 4.
+Let $\mathbb{E}$ be a differential field extension of $\mathbb{K}$ such that $\mathbb{L}=\mathbb{E}(e^{q_{1}(y^{1/\ell})},\ldots,e^{q_{n}(y^{1/\ell})})$.
+Then $\mathscr{T}=\mathrm{Gal}(\mathbb{L}/\mathbb{E})$ is called the exponential torus of (A.1).
+Thus, $\mathscr{T}$ is the group of all $\mathbb{E}$-automorphisms of $\mathbb{L}$.
+We next state the definition of the Stokes matrix for (4.2).
+Some preparations are needed to this end.
+Let $q(z)\in z^{-1}\mathbb{C}[z^{-1}]$, so that $q(z^{-1})$ is a polynomial of $z$.
+Let $k$ be the degree of $q(z^{-1})$ and let $q(z)=c_{0}z^{-k}+\ldots$, where $c_{0}\in\mathbb{C}$ is a constant.
+The direction $d$ determined by $\mathrm{Re}\,(c_{0}y^{-k/\ell})=0$, in which $e^{q(z)}$ changes its behavior, is called the Stokes direction for $q(z)$.
+Let $d_{1},d_{2}$ be consecutive Stokes directions.
+We say that $(d_{1},d_{2})$ is a negative Stokes pair if $\mathrm{Re}\,(c_{0}y^{-k/\ell})<0$ for $\arg(y)\in(d_{1},d_{2})$.
+We refer to the bisector a negative Stokes pair as a singular direction of $q(z)$.
+Figure 3. Singular direction $d$ and the two directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$.
+We now consider the formal fundamental matrix (A.2).
+Let $d$ be a singular direction of $q_{j}(z)-q_{l}(z)$ for some $j,l=1,\ldots,n$ with $j\neq l$.
+Choose two directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$ counterclockwise nearby such that the associated sectors overlap and contain $d$.
+See Fig.
+3.
+Let $\hat{Y}_{\mathrm{L}}(y)$ and $\hat{Y}_{\mathrm{R}}(y)$ be the multisums of $\hat{Y}(y)$ in the directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$, respectively.
+Then there exists $S_{jl}\in\mathrm{GL}(n;\mathbb{C})$ such that
+$\hat{Y}_{\mathrm{L}}(y)y^{L}e^{Q(1/z)}=\hat{Y}_{\mathrm{R}}(y)y^{L}e^{Q(1/z)}S_{jl}$ (A.3)
+since Eq.
+(A.2) gives a formal fundamental matrix for both cases.
+The matrix $S_{jl}$ is independent of the choice of $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$ and called a Stokes matrix for the singular direction $d$.
+We have the following on the Stokes matrices (see, e.g., Proposition 2.37 of [21] for its proofs).
+Proposition A.1.
+The Stokes matrices are unipotent such that all the diagonal elements are one and the $(j^{\prime},l^{\prime})$-element is zero if $d$ is a singular direction for $q_{j^{\prime}}(z)-q_{l^{\prime}}(z)$.
+Finally, we state one of the most important results on local differential Galois groups near irregular singularities as follows.
+Theorem A.2 (Ramis [24, 25]).
+The local differential Galois group of (4.2) around the irregular singular point $y=0$ is the Zariski closure of the group generated the formal monodromy $\hat{M}$, the exponential torus $\mathscr{T}$ and the Stokes matrices $S_{jl}$, $j,l=1,\ldots,n$ with $j\neq l$.
+See also [15, 23] for a proof of Theorem A.2. Theorem A.2 is regarded as a generalization of Schlesinger’s theorem (see, e.g., [23, 29]) for regular singularities to irregular ones.
+Appendix B Kovacic’s Result
 In this appendix we review necessary parts of Kovacic [13] on second-order linear differential equations on $\mathbb{C}(x)$.
-A.1. Fundamental results
+B.1. Fundamental results
 Consider second-order linear differential equations of the form
-$v_{xx}=r(x)v,$ (A.1)
+$v_{xx}=r(x)v,$ (B.1)
 where $r(x)\in\mathbb{C}(x)\setminus\mathbb{C}$.
-Let $G$ be the differential Galois group of (A.1) over $\mathbb{C}(x)$.
+Let $G$ be the differential Galois group of (B.1) over $\mathbb{C}(x)$.
 We easily see that $G$ is an algebraic subgroup of $\mathrm{SL}(2,\mathbb{C})$ (see, e.g., Chapter 6 of [12]).
-We have the following four cases for (A.1) (see [10, 13] for a proof).
-Proposition A.1 (Kaplansky [12], Kovacic[13]).
+We have the following four cases for (B.1) (see [10, 13] for a proof).
+Proposition B.1 (Kaplansky [12], Kovacic[13]).
 One of the following four cases occurs::\,::
@@
     $G=\mathrm{SL}(2,\mathbb{C})$.
-If one of cases (a)-(c) of Proposition A.1 occurs, then the connected identity component $G^{0}$ of $G$ is solvable and Eq.
-(A.1) is solvable.
+If one of cases (a)-(c) of Proposition B.1 occurs, then the connected identity component $G^{0}$ of $G$ is solvable and Eq.
+(B.1) is solvable.
 On the other hand, if cases (a)-(c) do not occur, then $G=G^{0}=\mathrm{SL}(2,\mathbb{C})$, i.e., Eq.
-(A.1) is not integrable in the sence of differential Galois theory.
+(B.1) is not integrable in the sence of differential Galois theory.
 Let $r(x)=r_{1}(x)/r_{2}(x)$, where $r_{1}(x),r_{2}(x)\in\mathbb{C}[z]$ are relatively prime.
@@
 By the order of $r(x)$ at $\infty$, we also mean the order of $\infty$ as a zero of $r(x)$, i.e., $\deg r_{2}(x)-\deg r_{1}(x)$.
-We have the following necessary conditions for cases (a)-(c) in Proposition A.1 to occur (see [10, 13] for a proof).
-Proposition A.2 (Kovacic [13]).
-The following conditions are necessary for the corresponding cases in Proposition A.1 to occur::\,::
+We have the following necessary conditions for cases (a)-(c) in Proposition B.1 to occur (see [10, 13] for a proof).
+Proposition B.2 (Kovacic [13]).
+The following conditions are necessary for the corresponding cases in Proposition B.1 to occur::\,::
   • (a)
@@
     No pole of $r(x)$ is of order greater than two and the order of $r(x)$ at $\infty$ is greater than one.
-Note that cases (a), (b) and (c) of Proposition A.1 may not occur even though conditions (a), (b) and (c) hold, respectively.
-In particular, Proposition A.2 only states fewer conditions for case (c) than the original version in [13] since they are enough for our application in Section 5.
-Kovacic [13] also gave an algorithm for finding a “closed-form” solution and detecting whether cases (a)-(c) in Proposition A.1 really occur or not under conditions (a), (b) and (c) of Proposition A.1, respectively.
+Note that cases (a), (b) and (c) of Proposition B.1 may not occur even though conditions (a), (b) and (c) hold, respectively.
+In particular, Proposition B.2 only states fewer conditions for case (c) than the original version in [13] since they are enough for our application in Section 5.
+Kovacic [13] also gave an algorithm for finding a “closed-form” solution and detecting whether cases (a)-(c) in Proposition B.1 really occur or not under conditions (a), (b) and (c) of Proposition B.1, respectively.
 In the following we describe necessary parts of his algorithm for cases (a) and (b) in the setting required in Section 5: The order of $r(x)$ at $\infty$ is zero.
 See [10, 13] for the full version of his algorithm and its proof.
-A.2. Case (a)
-We assume that condition (a) in Proposition A.2 holds as well as the order of $r(x)$ at $\infty$ is zero. When $g(x)$ has a pole of order $\ell\geq 2$ at $x=s$ and is expanded in a Laurent series
+B.2. Case (a)
+We assume that condition (a) in Proposition B.2 holds as well as the order of $r(x)$ at $\infty$ is zero. When $g(x)$ has a pole of order $\ell\geq 2$ at $x=s$ and is expanded in a Laurent series
 $g(x)=\sum_{j\geq-\ell}g_{j}(x-s)^{j}$
@@
 For each sequence $c=\{c(s)\}_{s\in\mathscr{S}\cup\{\infty\}}$ with $c(s)=+$ or --, define
-$d_{c}=\kappa_{\infty}^{c(\infty)}-\sum_{s\in\mathscr{S}}\kappa_{s}^{c(s)}.$ (A.2)
-If $d_{c}$ is not a non-negative integer for any sequence $c$, then case (a) of Proposition A.1 does not occur.
+$d_{c}=\kappa_{\infty}^{c(\infty)}-\sum_{s\in\mathscr{S}}\kappa_{s}^{c(s)}.$ (B.2)
+If $d_{c}$ is not a non-negative integer for any sequence $c$, then case (a) of Proposition B.1 does not occur.
 We also have a further recipe to determine whether it occurs or not when $d_{c}$ is a non-negative integer, although it is not given here since it is not necessary for our purpose.
 The reader who is interested in it should consult [10, 13].
-A.3. Case (b)
-We assume that condition (b) in Proposition A.2 holds as well as the order of $r(x)$ at $\infty$ is zero. For $s\in\mathscr{S}\cup\{\infty\}$ we define a set $E_{s}\subset\mathbb{Z}$ as follows:
+B.3. Case (b)
+We assume that condition (b) in Proposition B.2 holds as well as the order of $r(x)$ at $\infty$ is zero. For $s\in\mathscr{S}\cup\{\infty\}$ we define a set $E_{s}\subset\mathbb{Z}$ as follows:
   1. ($s_{1}$)
@@
 For each tuple $e=\{e_{s}\in E_{s}\mid s\in\mathscr{S}\}$, let
-$d_{e}=-\tfrac{1}{2}\sum_{s\in\mathscr{S}}e_{s}.$ (A.3)
-If $d_{e}$ is not a non-negative integer for any $e$, then case (b) of Proposition A.1 does not occur.
+$d_{e}=-\tfrac{1}{2}\sum_{s\in\mathscr{S}}e_{s}.$ (B.3)
+If $d_{e}$ is not a non-negative integer for any $e$, then case (b) of Proposition B.1 does not occur.
 Let $E=\{e\mid d_{e}\in\mathbb{N}_{0}\}$ and let
-$\theta(x)=\tfrac{1}{2}\sum_{s\in\mathscr{S}}\frac{e_{s}}{x-s}$ (A.4)
-for $e\in E$. By (A.3) and (A.4) we see that the Laurent series of $\theta(x)$ at $\infty$ is given by
+$\theta(x)=\tfrac{1}{2}\sum_{s\in\mathscr{S}}\frac{e_{s}}{x-s}$ (B.4)
+for $e\in E$. By (B.3) and (B.4) we see that the Laurent series of $\theta(x)$ at $\infty$ is given by
 $\theta(x)=-\frac{d_{e}}{x}+\cdots.$
@@
 $P^{\prime\prime\prime}+3\theta(x)P^{\prime\prime}+(3\theta(x)^{2}+3\theta^{\prime}(x)-4r(x))P^{\prime}$
-$+(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}-4r(x)\theta(x)-2r^{\prime}(x))P=0,$ (A.5)
-then case (b) of Proposition A.1 occurs and
-$v(x)=\exp\left(\int\tfrac{1}{2}\left(-\hat{\theta}(x)\pm\sqrt{4r(x)-\hat{\theta}(x)^{2}-2\hat{\theta}^{\prime}(x)}\right)\mathrm{d}x\right)$ (A.6)
-are solutions to (A.1), where $\hat{\theta}(x)=\theta(x)+P^{\prime}(x)/P(x)$. If such a polynomial is not found for any sequence $e$, then it does not occur.
+$+(\theta^{\prime\prime}(x)+3\theta(x)\theta^{\prime}(x)+\theta(x)^{3}-4r(x)\theta(x)-2r^{\prime}(x))P=0,$ (B.5)
+then case (b) of Proposition B.1 occurs and
+$v(x)=\exp\left(\int\tfrac{1}{2}\left(-\hat{\theta}(x)\pm\sqrt{4r(x)-\hat{\theta}(x)^{2}-2\hat{\theta}^{\prime}(x)}\right)\mathrm{d}x\right)$ (B.6)
+are solutions to (B.1), where $\hat{\theta}(x)=\theta(x)+P^{\prime}(x)/P(x)$. If such a polynomial is not found for any sequence $e$, then it does not occur.